A plane is cruising at an altitude of 24,000 ft. The plane begins to descend at a rate of 20 ft per sec. What equation gives the altitude h (in feet) of the plane as a function of the time t (in seconds) since it began descending? What is the altitude of the plane after 5 minutes?
Answer by mananth(16946)

h the height at any time after descent = 24000-20s
where s is in second
h= 24000-20*5*60 ( 1min = 60 sec)
h= 24000-6000
h= 18000 ft